# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ 3.3.1 Vertex extremal length and macroscopic circles
Uniform boundedness on compact sets. For $z\neq 0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity. Let $\gamma_{K}$ be a deterministic curve formed by the union of the line segment from 0 to $(2d_{K}+3)i$ and a curve from $(2d_{K}+3)i$ to $2+(d_{K}+2)i$, such that it is disjoint from $B(0;d_{K}+1)$ and separates $B(0;{d_{K}+1})$ from infinity. Then for any $\delta>0$, with positive probability, the planar Brownian motion started from 0 stays within the $\delta/2$-neighborhood of $\gamma_{K}$ and form a loop around $K\cup\{1\}\subset B(0;d_{K}+1)$ during time $[0,2]$. In particular, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}\leq\tau_{1}]$ is uniformly positive for $z\in K$ and small enough $\varepsilon$. Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, for each $z\in K$, there exists $\varepsilon^{\prime}>0$ such that $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}]>3\varepsilon^{\prime}$ for $\varepsilon$ sufficiently small. On the other hand, for each $\varepsilon^{\prime}>0$, there exists $M,\delta_{1}>0$, such that the probability that a Brownian motion from 0 separates $B(1;\delta_{1})$ from infinity before leaving $B(0;M)$ is at least $1-\varepsilon^{\prime}$. Using the convergence in law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ to standard Brownian motion, for $\varepsilon$ sufficiently small, with probability at least $1-2\varepsilon^{\prime}$, the trace of $\widehat{X}^{0,\rm circle,\widetilde{\varepsilon}}$ separates 1 from infinity before leaving $B(0;2M)$. In other words, for $\varepsilon$ sufficiently small, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}\leq\tau_{1}]\leq 2\varepsilon^{\prime}$ for $|z|>2M$. Thus $|f^{\varepsilon}(z)|\leq 2M$ for all $z\in K$ and small enough $\varepsilon$, which gives the uniform boundedness for $f^{\varepsilon}$ on compacts.
@@
Equicontinuity near 0 and 1. Let $\delta>0$. For $z\neq 0,1$, we work on the event $\{\tau_{z}=\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity at the same time. Using the same $\gamma_{K}$ and Brownian motion argument as in the previous step, one can see that there exists $\delta^{\prime}>0$ such that for $\varepsilon$ small enough $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}^{0}[\tau_{z}=\tau_{1}]>\delta^{\prime}$ when $z\in K$ and $|z|>\delta$. On the other hand, there exists $\delta^{\prime\prime}>0$ such that the probability that a planar Brownian motion separates $B(0;2\delta^{\prime\prime})$ from infinity before leaving $B(0;1/2)$ is at least $1-\delta^{\prime}/2$, and thus using the convergence in law of $\widehat{X}^{\varepsilon}$ to Brownian motion in local uniform topology, for $|z|<\delta^{\prime\prime}$ and all sufficiently small $\varepsilon$, $\mathbb{P}_{\varepsilon}^{0}[\tau_{z}=\tau_{1}]<\delta^{\prime}$. Since $f^{\varepsilon}(\widehat{X}^{\varepsilon})$ and $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ are coupled to have the same trace, we conclude that there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)|<\delta$. This verifies the equicontinuity near 0. The equicontinuity near 1 can be verified similarly by considering walks started at 1.
@@
Equicontinuity at other points in $K$. Fix $z\in K$ with $z\neq 0,1$. For $z^{\prime}\neq z,0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}\}$, i.e., when the path separates 1 from infinity, it separates $z$ from infinity but does not separate $z^{\prime}$ from infinity. From the previous cases, $\{f^{\varepsilon}(z)\}$ is uniformly bounded away from 0, 1 and $\infty$. Fix $0<\delta<\frac{1}{2}\min_{\varepsilon>0}\{1,|f^{\varepsilon}(z)|,|f^{\varepsilon}(z)-1|\}$. Let $\gamma_{f^{\varepsilon}(z)}$ be a finite path starting from 0 that first separates $B(f^{\varepsilon}(z);\delta/4)$ and then separates $B(1;\delta/4)$ from infinity, but does not separate any points in $\mathbb{C}\backslash(B(f^{\varepsilon}(z);{\delta/2})\cup B(1;\delta/2))$ from infinity. Then using a similar argument in previous cases, where the Brownian motion has positive probability to stay sufficiently close to $\gamma_{f^{\varepsilon}(z)}$, and the random walks converges in law to Brownian motion in local uniform topology, we see that for each $\delta>0$ there exists $\delta^{\prime},\delta^{\prime\prime}>0$ such that for sufficiently small $\varepsilon$, $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}[\tau_{f^{\varepsilon}(z)}\leq\tau_{1}<\tau_{z^{\prime}}]>\delta^{\prime}$ if $|z^{\prime}-f^{\varepsilon}(z)|>\delta$ whereas $\mathbb{P}_{\varepsilon}[\tau_{z}\leq\tau_{1}<\tau_{z^{\prime}}]<\delta^{\prime}$ if $|z^{\prime}-z|<\delta^{\prime\prime}$. In particular, there exists $\varepsilon_{0}$, such that for any $\varepsilon<\varepsilon_{0}$ and $|z-z^{\prime}|<\delta^{\prime\prime}$, $|f^{\varepsilon}(z)-f^{\varepsilon}(z^{\prime})|<\delta$.
@@
This finishes the proof for the tightness of $\{f^{\varepsilon}\}|_{z\in K}$ for each compact set $K$. By an identical argument, $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight for each compact set $K$. By the Arzelà-Ascoli theorem, for any subsequence of $\varepsilon$, there exists a further subsequence such that $f^{\varepsilon}$ converges uniformly on compacts to a continuous function $f$ on $\mathbb{C}$ such that $f(0)=0,f(1)=1$. By using that $\{(f^{-1})^{\varepsilon}|_{K}\}$ is also tight and taking a further subsequence we may assume that $f$ is a homeomorphism. Then if $(W_{t})_{t\geq 0}$ is a Brownian motion from $z$, $f(W_{t})_{t\geq 0}$ has the law of a planar Brownian motion from $f(z)$ modulo time change. Then by Proposition 3.14, $f$ must be either the identity function or the reflection. This completes the proof. ∎
@@ 4.1 Regularity estimates for the Riemann surface
In this section, we prove several regularity estimates for the Riemann surface $M(\mathcal{H})$ under a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ that will be applied in the proof of Theorem 1.8. In Lemmas 4.34.4, and 4.5, we prove that if we have a microscopic set in $\mathcal{H}$ (resp. $\varphi(M(\mathcal{H}))$), then the corresponding set in $\varphi(M(\mathcal{H}))$ (resp. $\mathcal{H}$) will also be microscopic. Let $D_{k}$ be the union of $\varphi(P_{H})$ of the cells $H\in\mathcal{H}$ intersecting $2^{k},2^{k}]^{2}$. In Lemma 4.6, we prove that the diameter of $D_{k}$ has a growth rate faster than $2^{k/l}$ for some deterministic $l$. In Lemma 4.8, we prove that for large $k$, $D_{k}$ contains a disk with diameter at least $c\mathrm{diam}(D_{k})$. These lead to Lemma 4.9, where we prove that $M(\mathcal{H})$ is a.s. parabolic. Finally in Lemma 4.10, we prove that the function mapping points in $\varphi(M(\mathcal{H}))$ to their corresponding cells in $\mathcal{H}$ has a polynomial growth, which will be applied in Section 4.4 in the proof of Theorem 1.8.
@@
For each square $S$, we define a simply connected surface $M(S)\subseteq M(\mathcal{H})$ as follows. For every $a>0$, let $\mathring{S}^{a}$ be the square concentric to $S$ with side length $a$. We work on the event that $\mathcal{H}(\mathring{S}^{a})$ is connected; otherwise we set $M(S;a)=\emptyset$. Let $M_{0}(S;a)\subseteq M(\mathcal{H})$ be the surface formed by the union of the (closed) equilateral triangles of $M(\mathcal{H})$ with at least one of the vertices in $\mathcal{H}(\mathring{S}^{a})$. Note that when viewed as a graph, $M_{0}(S;a)$ is 2-connected since $\mathcal{H}(\mathring{S}^{a})$ is connected. In fact, one can verify that, for any infinite plane triangulation $\mathcal{T}$ and a connected subgraph $\mathcal{G}_{1}\subseteq\mathcal{G}$, if we let $\mathcal{G}_{2}$ be the subgraph $\mathcal{G}$ formed by all the vertices in $\mathcal{G}_{1}$ and their neighbors, as well as all the edges connecting these vertices, then $\mathcal{G}_{2}$ is 2-connected. We further set $M(S;a)$ to be the union of $M_{0}(S;a)$ and the bounded connected components of $M(\mathcal{H})\backslash M_{0}(S;a)$. Then $M(S;a)$ has disk topology. Let $M(S)$ be the largest surface of the form $M(S;a)$ for some $a\in(0,|S|)$ such that for each vertex on $M(S;a)$, its cell is contained within $S$. Note that these objects are well defined since $\mathcal{H}$ is locally finite. See Figure 5 for an illustration.
Lemma 4.1.
Almost surely, for each given $\varepsilon>0$ and $a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $M(S;(1-\varepsilon)|S|)\subseteq M(S)$.
Proof.
By the connectivity in Definition 1.4, for each $b\in(0,1)$, almost surely for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $\mathcal{H}(\mathring{S}^{b|S|})$ is a connected subgraph of $\mathcal{H}$. Let $H\in\mathcal{H}$ be such that its vertex $v_{H}$ on $M(\mathcal{H})$ is in $M(S;(1-\varepsilon)|S|)$. Then there are two cases: (i) $H$ is adjacent to some vertex $H^{\prime}\in\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$ or (ii) there exists a loop consisting of vertices for $H_{1},...,H_{n}$ on $M(\mathcal{H})$ surrounding $v_{H}$, such that each $H_{j}$ are adjacent to some cell in $\mathcal{H}(\mathring{S}^{(1-\varepsilon)|S|})$. In the setting (i), by Lemma 2.6, if we assume that $k$ is large, then $H\subseteq S$. For the case (ii), by Lemma 2.6, for large $k$, $H_{1},...,H_{n}\in\mathcal{H}(\mathring{S}^{(1-3\varepsilon/4)|S|})$. Letting $\{z_{H}:H\in\mathcal{H}\}$ and $\{e_{H,H^{\prime}}:H\sim H^{\prime}\}$ be the planar embedding of the map $\mathcal{H}$ as in the almost planarity in Definition 1.5, for large enough $k$, $e_{H_{1},H_{2}},...,e_{H_{n},H_{1}}\subseteq\mathring{S}^{(1-\varepsilon/2)|S|}$ and thus $z_{H}\in\mathring{S}^{(1-\varepsilon/2)|S|}$. By the same condition and Lemma 2.6 it follows that $H\subseteq S$. This finishes the proof. ∎
@@
For each $k>0$, let $\varphi_{k}$ be the uniformization map sending $M(S_{k})$ to $|S_{k}|\mathbb{D}$ such that the vertex on $M(S_{k})$ for the cell containing the center of $S_{k}$ is sent to 0. Recall the definition of the semiflower in Lemma 4.2. For each $H\in\mathcal{H}$, we write $P_{H}$ for the semiflower for the vertex $H$ on the surface $M(\mathcal{H})$.
@@
Figure 7: An illustration of the proof of Lemma 4.3, where the left panel is a part of our cell configuration and the right panel is the corresponding image of the part of the surface $M(\mathcal{H})$ under a conformal map. The squares concentric to $S$ of side lengths $r_{0}$ and $r_{1}$ are colored in blue on the left panel. In (4.12), we prove that one can find some $\rho\in[r_{0},r_{1}]$, such that under a conformal map the union of the semi-flowers for the cells intersecting the square concentric to $S$ of side length $\rho$ (shaded in light blue on the right panel) has diameter less than $c(-\log\delta)^{-1/2}|S_{k}|$. Further using the line connectivity property and almost planarity in Definition 1.4 and Definition 1.5 gives the claim.
@@
If $\mathring{S}^{\rho}\subseteq\mathring{S}_{k}^{7|S_{k}|/8}$, using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ intersecting $\partial\mathring{S}^{\rho}$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells, and in the planar embedding $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ in Condition 4. form a loop around $\mathring{S}^{1.5\delta|S_{k}|}$, and thus encloses all the points $z_{H}$ for $H\in{S}$. This implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ is separated from infinity by $\cup_{j=1}^{n}\varphi(P_{H_{j}})$, which together with (4.12) implies (4.5).
If $\mathring{S}^{\rho}\nsubseteq\mathring{S}_{k}^{7{|S_{k}|}/8}$, since $\rho<r_{1}\leq|S_{k}|/4$, $\mathring{S}^{\rho}$ is disjoint from the cell with closest distance to the center of $S_{k}$. Using the conditions from Definition 1.4 and 1.5, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ either intersecting $\partial\mathring{S}^{\rho}$ or with their vertex on $M(\mathcal{H})$ on the boundary of the surface $M(S_{k})$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells and the edges $\gamma_{e_{1}},...,\gamma_{e_{n}}$ enclose a loop around all the points $z_{H}$ where $H\in\mathcal{H}(S)$ and the vertex for $H$ is on $M(S_{k})$. This together with (4.12) implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ can be separated from 0 by $\partial(|S_{k}|\mathbb{D})$ and some curves of total length $4c_{2}(\log\delta)^{-1/2}|S_{k}|$. This finishes the proof of (4.5).
@@
In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L^{*}(\rho)\leq 4c_{2}(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|$. Recall the points $\{z_{H}:H\in\mathcal{H}\}$ and the curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in the planar embedding from the almost planarity condition in Definition 1.5. Note that by our choice of $r_{1}$, $B(z;\rho)\subseteq|S_{k}|\mathbb{D}$. One can find edges $e_{1},...,e_{n}$ forming a loop in $\mathcal{H}$ such that $\gamma_{e_{1}},...,\gamma_{e_{n}}$ make a loop surrounding all the vertices $z_{H}$ with $\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset$, and the union of theses edges has diameter less than $20c_{2}(-\log\delta)^{-1/2}|S_{k}|$. Applying  (1.2) once more gives (4.13). ∎
We will also need the version of Lemmas 4.3 and 4.4 for $2^{k},2^{k}]^{2}$ instead of $S_{k}$. With slight abuse of notation, we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$. For each $k>0$, let $\widehat{\varphi}_{k}$ be the uniformization map sending $M([-2^{k},2^{k}]^{2})\subseteq M(\mathcal{H})$ to $2^{k}\mathbb{D}$ and sending $0\in M(\mathcal{H})$ to 0.

### 06 May 2026 v2

@@ 3.3.1 Vertex extremal length and macroscopic circles
Step 1: Uniform boundedness on $K$. For $z\neq 0,1$, we work on the event $\{\tau_{z}\leq\tau_{1}<\sigma_{\ell}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity and has not exited $B(0;2^{\ell-1}$). Let $\gamma_{K}$ be a deterministic curve formed by the union of the line segment from 0 to $(2d_{K}+3)i$ and a curve from $(2d_{K}+3)i$ to $2+(d_{K}+2)i$, such that it is disjoint from $B(0;d_{K}+1)$ and separates $B(0;{d_{K}+1})$ from infinity. Then for any $\delta>0$, with positive probability, the planar Brownian motion started from 0 stays within the $\delta/2$-neighborhood of $\gamma_{K}$ and form a loop around $K\cup\{1\}\subset B(0;d_{K}+1)$ during time $[0,2]$. In particular, a.s. for fixed $\ell$ and large enough $k$, $\mathbb{P}_{k,\ell}^{0}\big[\{\tau_{z}\leq\tau_{1}\}\cap\{\sup_{0\leq t\leq\tau_{1}}|\widehat{X}_{t}|\leq 2^{\ell-1}\}\big]$ is uniformly positive for $z\in K$. Since $f^{\varepsilon}(\widehat{X}^{k,\ell})$ and $\widehat{X}^{\mathrm{circle},k,\ell}$ are coupled to have the same trace until leaving $f^{k,\ell}([-2^{\ell-1},2^{\ell-1}]^{2})$ for each $z\in K$, there exists $\varepsilon^{\prime}>0$ such that $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}\leq\tau_{1}]>3\varepsilon^{\prime}$ for $k$ sufficiently large. On the other hand, for each $\varepsilon^{\prime}>0$, there exists $M,\delta_{1}>0$, such that the probability that a Brownian motion from 0 separates $B(1;\delta_{1})$ from infinity before leaving $B(0;M)$ is at least $1-\varepsilon^{\prime}$. Using the convergence in law of $\widehat{X}^{\mathrm{circle},k,\ell}$ to standard Brownian motion, for $k$ sufficiently large, with probability at least $1-2\varepsilon^{\prime}$, the trace of $\widehat{X}^{0,\mathrm{circle},k,\ell}$ separates 1 from infinity before leaving $B(0;2M)$. In other words, for $k$ sufficiently large, $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}\leq\tau_{1}]\leq 2\varepsilon^{\prime}$ for $|z|>2M$. Thus $|f^{k,\ell}(z)|\leq 2M$ for all $z\in K$, large enough $k$, and $\ell>\widetilde{\ell}+1$.
By a similar argument, for each compact set $K^{\prime}\subseteq\mathbb{C}$, there exists $\varepsilon^{\prime\prime}>0$ such that for every $\ell$ and large enough $k$, $\inf_{z^{\prime}\in K^{\prime}}\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z^{\prime}}\leq\tau_{1}]>\varepsilon^{\prime\prime}$, and for $\ell$ large enough and every $k$ large enough, $\mathbb{P}_{k,\ell}^{0}[\tau_{1}<\sigma_{\ell-8}]>1-\varepsilon^{\prime\prime}$. The latter inequality implies that for large enough $\ell$ and $k$, $\inf_{z^{\prime}\notin f^{k,\ell}(K)}\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z^{\prime}}\leq\tau_{1}]>1-\varepsilon^{\prime\prime}$, which further implies that almost surely, there exists some $\ell_{0}$, such that for all $\ell>\ell_{0}$, there is a $k_{\ell}>0$ such that $K^{\prime}\subseteq f^{k,\ell}(K)$ for all $k>k_{\ell}$.
@@
Step 2: Equicontinuity near 0. Let $\delta>0$. For $z\neq 0,1$, we work on the event $\{\tau_{z}=\tau_{1}\}$, i.e., when the trace separates 1 from infinity, it separates $z$ from infinity at the same time. Using the same $\gamma_{K}$ and Brownian motion argument as in the previous step, one can see that there exists $\delta^{\prime}>0$ such that for $k$ large enough $\mathbb{P}_{k,\ell}^{0}[\tau_{z}=\tau_{1}<\sigma_{\ell}]>\delta^{\prime}$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}=\tau_{1}]>\delta^{\prime}$ when $z\in K$ and $|z|>\delta$. On the other hand, there exists $\delta^{\prime\prime}>0$ such that the probability that a planar Brownian motion separates $B(0;2\delta^{\prime\prime})$ from infinity before leaving $B(0;1/2)$ is at least $1-\delta^{\prime}/2$, and thus using the convergence in law of $\widehat{X}^{\varepsilon}$ to Brownian motion in local uniform topology, for $|z|<\delta^{\prime\prime}$ and all sufficiently large $k$, $\mathbb{P}_{k,\ell}^{0}[\tau_{z}<\tau_{1};\tau_{z}<\sigma_{\ell}]>1-\delta^{\prime}$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{z}<\tau_{1}]>1-\delta^{\prime}$. Since $f^{k,\ell}(\widehat{X}^{k,\ell})$ and $\widehat{X}^{\mathrm{circle},k,\ell}$ are coupled to have the same trace until leaving $f^{k,\ell}(K)$, we have a.s. for large $k$, $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}=\tau_{1}]>\delta^{\prime}$ for $|z|>\delta$ and $\widetilde{\mathbb{P}}_{k,\ell}^{0}[\tau_{f^{k,\ell}(z)}<\tau_{1}]>1-\delta^{\prime}$ for $|z|<\delta^{\prime\prime}$. We conclude that there a.s. exists $k_{0}$, such that for any $k>k_{0}$, if $|z|<\delta^{\prime\prime}$ then $|f^{k,\ell}(z)|<\delta$, and if $|z|\geq\delta$ then $|f^{k,\ell}(z)|\geq\delta^{\prime\prime}$. This verifies the equicontinuity near 0.
@@
Step 3: Equicontinuity at other points in $K$. This step is almost identical to the argument in step 2 by considering walks started from a fixed $z\in K$ with $z\neq 0$. From the previous cases, for fixed $\ell$, uniformly in $k$, $\{f^{k,\ell}(z)\}$ is bounded away from 0 and $\infty$. Pick $0<\delta_{z}<\min\{|z|/2,1\}$, fix $0<\delta<\frac{1}{2}\min\{|f^{k,\ell}(w)|:k\geq 10,|w-z|<\delta_{z}\}$, $M\geq\sup_{k>0}|f^{k,\ell}(K)|$ and $z^{\prime\prime}\in B(0;2M)\backslash B(0;2\delta)$ with $|z^{\prime\prime}-f^{k,\ell}(z)|>\delta$. Let $\gamma_{f^{k,\ell}(z),z^{\prime\prime}}$ be a finite path starting from $f^{k,\ell}(z)$ that first separates $B(0;\delta/4)$ and then $B(z^{\prime\prime};\delta/4)$ from infinity, but does not separate any points in $\mathbb{C}\backslash(B(z^{\prime\prime};{\delta/2})\cup B(0;\delta/2))$ from infinity. Then using a similar argument in previous cases, where the Brownian motion has positive probability to stay sufficiently close to $\gamma_{f^{k,\ell}(z),z^{\prime\prime}}$, and the random walks converges in law to Brownian motion in local uniform topology, we see that for each $\delta>0$ there exists $\delta^{\prime}>0$ such that a.s. for sufficiently large $k$, $\widetilde{\mathbb{P}}_{k,\ell}^{f^{k,\ell}(z)}[\tau_{0}<\tau_{z^{\prime\prime}}]>\delta^{\prime}$ if $|z^{\prime\prime}-f^{k,\ell}(z)|>\delta$ and $2\delta<|z^{\prime\prime}|<2M$. On the other hand, for some $\delta^{\prime\prime}>0$, for all $k$ large enough, $\mathbb{P}_{k,\ell}^{z}[\tau_{z^{\prime}}<\tau_{0};\tau_{z^{\prime}}<\sigma_{\ell}]>1-\delta^{\prime}$ for all $|z^{\prime}-z|<\delta^{\prime\prime}$, which further implies that $\widetilde{\mathbb{P}}_{k,\ell}^{f^{k,\ell}(z)}[\tau_{f^{k,\ell}(z^{\prime})}<\tau_{0}]>1-\delta^{\prime}$. This implies a.s. for large enough $k$, $|f^{k,\ell}(z^{\prime})-f^{k,\ell}(z)|<\delta$ whenever $|z-z^{\prime}|\leq\min\{\delta^{\prime\prime},\delta_{z}\}$ and $z^{\prime}\in K$.
@@
This finishes the proof for the tightness of $\{\{f^{k,\ell}\}|_{z\in K}:k\in\mathbb{N},\ell>\widetilde{\ell}+1\}$ for each $\widetilde{\ell}>10$ and $K=[-2^{\widetilde{\ell}-5},2^{\widetilde{\ell}-5}]^{2}$. By the Arzelà-Ascoli theorem, for any subsequence of $k$, there exists a further subsequence $k_{n},\ell_{n}\to\infty,$ such that for every $\ell>10$, $f^{k_{n},\ell_{n}}$ converges uniformly on $2^{\ell-5},2^{\ell-5}]^{2}$ to a continuous function $f$ on $\mathbb{C}$ such that $f(0)=0,f(1)=1$. By the last sentence of Step 1, we have $f(\mathbb{C})=\mathbb{C}$. Using an identical argument as Step 2 and Step 3, for each compact set $K^{\prime}$, for each large enough $\ell$, $\{(f^{k,\ell})^{-1}|_{K^{\prime}}\}$ is also tight. By taking a further subsequence we may assume that $f$ is a homeomorphism. Then if $(W_{t})_{t\geq 0}$ is a Brownian motion from $z$, $f(W_{t})_{t\geq 0}$ has the law of a planar Brownian motion from $f(z)$ modulo time change. Then by Proposition 3.12, $f$ must be either the identity function or the reflection $f(z)=\overline{z}$. To rule out the $f(z)=\overline{z}$ case, consider cells $H_{1},...,H_{m}$ in $\mathcal{H}{(\partial[-2^{k},2^{k}]^{2})}$ such that the edges $\gamma_{e}^{k+1}$ connecting the points $z_{H_{j}}^{k+1}$ form a clockwise loop which is close to $\partial[-2^{k},2^{k}]^{2}$. Then the corresponding loop formed by the disks for $H_{1},...,H_{m}$ is also clockwise, which implies that we cannot have $f(z)=\overline{z}$. ∎
@@ 4.1 Regularity estimates for the Riemann surface
In this section, we prove several regularity estimates for the Riemann surface $M(\mathcal{H})$ under a uniformization map $\varphi:M(\mathcal{H})\to\mathbb{C}$ that will be applied in the proof of Theorem 1.7. In Lemmas 4.24.3, and 4.4, we prove that if we have a microscopic set in $\mathcal{H}$ (resp. $\varphi(M(\mathcal{H}))$), then the corresponding set in $\varphi(M(\mathcal{H}))$ (resp. $\mathcal{H}$) will also be microscopic. Let $D_{k}$ be the union of $\varphi(P_{H})$ of the cells $H\in\mathcal{H}$ intersecting $2^{k},2^{k}]^{2}$. In Lemma 4.5, we prove that the diameter of $D_{k}$ has a growth rate faster than $2^{k/l}$ for some deterministic $l$. In Lemma 4.7, we prove that for large $k$, $D_{k}$ contains a disk with diameter at least $c\mathrm{diam}(D_{k})$. These lead to Lemma 4.8, where we prove that $M(\mathcal{H})$ is a.s. parabolic. Finally in Lemma 4.9, we prove that the function mapping points in $\varphi(M(\mathcal{H}))$ to their corresponding cells in $\mathcal{H}$ has a polynomial growth, which will be applied in Section 4.4 in the proof of Theorem 1.7.
@@
For a square $S$, recall the disk triangulation $\mathcal{M}(S)\subseteq\mathcal{M}$ defined in Lemma 2.8, and the Riemann surface $M(S)$ associated with $\mathcal{M}(S)$. For each $k>0$, let $\varphi_{k}$ be the uniformization map sending $M(S_{k})$ to $|S_{k}|\mathbb{D}$ such that the vertex on $M(S_{k})$ for the cell containing the center of $S_{k}$ is sent to 0. Recall the definition of the semiflower in Lemma 4.1. For each $H\in\mathcal{H}$, we write $P_{H}$ for the semiflower for the vertex $H$ on the surface $M(\mathcal{H})$.
@@
Figure 8: An illustration of the proof of Lemma 4.2, where the left panel is a part of our cell configuration and the right panel is the corresponding image of the part of the surface $M(\mathcal{H})$ under a conformal map. The squares concentric to $S$ of side lengths $r_{0}$ and $r_{1}$ are colored in blue on the left panel. In (4.12), we prove that one can find some $\rho\in[r_{0},r_{1}]$, such that under a conformal map the union of the semi-flowers for the cells intersecting the square concentric to $S$ of side length $\rho$ (shaded in light blue on the right panel) has diameter less than $c(-\log\delta)^{-1/2}|S_{k}|$. Further using the line connectivity property and almost planarity condition from Definition 1.4 and Proposition 2.15 gives the claim.
@@
Now consider the embedding $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ in Proposition 2.15. If $\mathring{S}^{\rho}\subseteq\mathring{S}_{k}^{7|S_{k}|/8}$, using the conditions from Definition 1.4 and Proposition 2.15, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ intersecting $\partial\mathring{S}^{\rho}$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells, and in the planar embedding $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ the edges $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ form a loop around $\mathring{S}^{1.5\delta|S_{k}|}$, and thus encloses all the points $z_{H}$ for $H\in{S}$. This implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ is separated from infinity by $\cup_{j=1}^{n}\varphi(P_{H_{j}})$, which together with (4.12) implies (4.5).
If $\mathring{S}^{\rho}\nsubseteq\mathring{S}_{k}^{7{|S_{k}|}/8}$, since $\rho<r_{1}\leq|S_{k}|/4$, $\mathring{S}^{\rho}$ is disjoint from the cell with closest distance to the center of $S_{k}$. Using Definition 1.4 and the embedding above, for large enough $k$, there exists cells $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ either intersecting $\partial\mathring{S}^{\rho}$ or with their vertex on $M(\mathcal{H})$ on the boundary of the surface $M(S_{k})$, such that there are edges $e_{1},...,e_{n}$ forming a loop with these cells and the edges $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ enclose a loop around all the points $z_{H}$ where $H\in\mathcal{H}(S)$ and the vertex for $H$ is on $M(S_{k})$. This together with (4.12) implies that $\{\varphi_{k}(P_{H}):H\in\mathcal{H}({S})\}$ can be separated from 0 by $\partial(|S_{k}|\mathbb{D})$ and some curves of total length $4c_{2}(\log\delta)^{-1/2}|S_{k}|$. This finishes the proof of (4.5).
@@
In particular, there exists $\rho\in[r_{0},r_{1}]$ such that $L^{*}(\rho)\leq 4c_{2}(\log(1-\lambda)-\log 16\delta)^{-1/2}|S_{k}|$. Recall the points $\{z_{H}^{k+3}:H\in\mathcal{V}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ and the curves $\{\gamma_{e}^{k+3}:e\in\mathcal{E}\mathcal{M}([-2^{k+3},2^{k+3}]^{2})\}$ in the planar embedding from Proposition 2.15. Note that by our choice of $r_{1}$, $B(z;\rho)\subseteq|S_{k}|\mathbb{D}$. One can find edges $e_{1},...,e_{n}$ forming a loop in $\mathcal{H}$ such that $\gamma_{e_{1}}^{k+3},...,\gamma_{e_{n}}^{k+3}$ make a loop surrounding all the vertices $z_{H}^{k+3}$ with $\varphi_{k}(P_{H})\cap B(z;r)\neq\emptyset$, and the union of theses edges has diameter less than $20c_{2}(-\log\delta)^{-1/2}|S_{k}|$. Applying  (2.12) once more gives (4.13). ∎
We will also need the version of Lemmas 4.2 and 4.3 for $2^{k},2^{k}]^{2}$ instead of $S_{k}$. With slight abuse of notation, we write 0 for the vertex on $M(\mathcal{H})$ representing the cell $H_{0}$. For each $k>0$, let $\widehat{\varphi}_{k}$ be the uniformization map sending $M([-2^{k},2^{k}]^{2})\subseteq M(\mathcal{H})$ to $2^{k}\mathbb{D}$ and sending $0\in M(\mathcal{H})$ to 0.